How many 5 mg doses of valium can the manufacturer produce from a 200 kg batch synthesis?
1
Expert's answer
2018-02-15T07:14:46-0500
5 mg=5*10^(-6)kg
200 kg=2*10^(2)kg
N= 2*10^(2)kg/ 5*10^(-6)kg=4*10^(7)=40 000 000
Answer: 40 000 000 doses